Five pencils and three erasers cost 65 cents. Three pencils and two erasers cost 40 cents. Find the cost of one pencil and one eraser.
Answer by checkley77(12844)

5P+3E=65 MULTIPLY THIS EQUATION BY -3
3P+2E=40 MULTIPLY THIS EQUATION BY 5 & ADD THEM.
-15P-9E=-195
15P+10E=200
-------------------------------
E=5 COST OF ONE ERASER.
5P+3*5=65
5P+15=65
5P=65-15
5P=50
P=10 COST OF ONE PENCIL.
PROOF
3*10+2*5=40
30+10=40
40=40